I have slightly rotated the lambda/2 plate, which is used for attenuating the REFL beam's power on the AS table
because the plate had been at an unusual angle for investigation of the glitches since last Thursday.
It means the laser power going to the coating thermal noise setup has also changed. Just keep it in mind.
So today we set up the Jenny RC temperature setup to lock the LWE NPRO to the RC and then set up the beat note with the IFO REFL beam on the AS table. By using the 2 laser beat, we are avoiding the VCO phase noise issue which used to limit the PSL frequency noise at ~0.01 Hz/rHz. To do this we have reworked some of the optics on the PSL and AS tables, but I think its been done without disturbing the beams for the regular locking. Beat note has been found, but the NPRO has still not been locked to the RC - next we setup the lockin amp, dither the PZT, and then use the New Focus lock box to lock it to the RC.